Smarte Plc is a recently listed company that manufactures and sells mobile phones. Its competitive advantage is that it manufactures and sells LazerFon, a new mobile phone with a built in laser projector. The table below presents essential information in order to compute the components needed to value Smarte using DCF valuation. Sales are reported in Emil. 2010 2011 2012 2013 2014 2015 2016 SS Sales Growth rate in sales 100 50.0% | 50.0% | 20.0% | 12.0% | 12.0% | 12.0% | 45% Depreciation10 1010 10 10 10 10 Cost of Sales-to- Sales Net Working Capital-to-Sales SG&A Expense- to-Sales CAPEX-to- Depreciation 60.60.6 0.6 0.6 0.6 0.6 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.04 | 0.041 0.04 | 0.04 | 0.04| O041004 4 4 4 4 4 Assume the change in net working capital between 2010 and 2011 was zero. The corporate tax rate is 25% and the projected equity risk premium is 4.5%. The risk free rate is also 4.5%. Smartes cost of debt is 5% and its debt to capital ratio is 40%, its beta is 1.1 and its market value of debt is 300m. Assume it is now January 2011. 36. The firms cost of equity is: (a) (b) (c) (d) 10.20% 9.45 7.17% None of the above 37. The firms NOPAT in 2015 and beyond 2016 is approximately (a) (b) (c) (d) 112mil and 133mil 84 mil and 100mil 53mil and 70mil None of the above CM108 2018/9 A 001 TURN OVER
Answer #1

(36) Equity Risk Premium = Cost of Equity - Risk-Free Rate = 4.5 %

Cost of Equity - 4.5 = 4.5

Cost of Equity = 9 %

Hence, the correct option is (d).

(37)

Sales in 2010 = 100

Sales in 2015 = 100 x 1.5 x 1.5 x 1.2 x 1.12 x 1.12 = 338.688

Cost of Sales - Sales = 0.6 and SG&A Expense - Sales = 0.04

Cost of Sales = 0.6 x 338.688 = 203.2128 and SG&A Expense = 338.688 x 0.04 = 13.54752

EBITDA = Sales - Cost of Sales - SG&A Expense = 338.688 - 203.2128 - 13.54572 = 121.9277

Less: Depreciation = 10

EBIT = 111.9277

Less: Tax @ 25 % = 0.25 x 111.9277 = 27.98192

NOPAT = 83.94576 ~ 84

Sales beyond 2016 = 100 x 1.5 x 1.5 x 1.2 x 1.12 x 1.12 x 1.12 x 1.045 = 396.4004

Cost of Sales - Sales = 0.6 and SG&A Expense - Sales = 0.04

Cost of Sales = 0.6 x 396.4004 = 237.8403 and SG&A Expense = 396.4004 x 0.04 = 15.85602

EBITDA = Sales - Cost of Sales - SG&A Expense = 396.4004 - 237.8403 - 15.85602  = 142.7042

Less: Depreciation = 10

EBIT = 132.7042

Less: Tax @ 25 % = 0.25 x 132.7042 = 33.17604

NOPAT = 99.52812 ~ 100

Hence, the correct option is (b).